# LionHeart: A Layer-based Mapping Framework for Heterogeneous Systems with Analog In-Memory Computing Tiles

### v1

@@ Abstract
When arranged in a crossbar configuration, resistive memory devices can be used to execute Matrix-Vector Multiplications, the most dominant operation of many Machine Learning (ML) algorithms, in constant time complexity. Nonetheless, when performing computations in the analog domain, novel challenges are introduced in terms of arithmetic precision and stochasticity, due to non-ideal circuit and device behaviour. Moreover, these non-idealities have a temporal dimension, resulting in a degrading application accuracy over time. Facing these challenges, we propose a novel framework, named LionHeart, to obtain hybrid analog-digital mappings to execute Deep Learning (DL) inference workloads using heterogeneous accelerators. The accuracy-constrained mappings derived by LionHeart showcase, across different Deep Neural Networks (DNNs) and datasets, high accuracy and potential for speedup. The results of the full system simulations highlight run-time reductions and energy efficiency gains that exceed 6$\times$, with a user-defined accuracy threshold with respect to a fully digital floating point implementation.
@@ I Introduction
Deep Learning-based solutions have been applied in various industrial and scientific applications, including computer vision, natural language processing, and speech recognition. Nowadays, State-Of-The-Art (SOTA) DL models require significant memory to store their parameters (i.e., weights and biases), and billions of Multiply-Accumulate (MAC) operations per inference. These extreme computing and memory requirements pose a challenge for deploying DL workloads on edge devices, whose form factor and energy budget require limited computing capabilities and memory size. However, the execution of DL workloads at the edge represents a very attractive prospect, as it would allow the processing of data samples close to where it is collected, resulting in several benefits, including reduced latency, a higher degree of security and privacy, improved bandwidth efficiency, as well as increased resiliency. Therefore, many research works have proposed strategies to optimize DL workloads for edge scenarios.
Figure 1: LionHeart considers the mapping of FC and CONV DNN layers to heterogeneous PUs featuring AIMC tiles and general-purpose cores to maximize performance while adhering to a user-specified accuracy constraint.
@@
ML algorithms, such as DNNs and transformers, are structured as regular sequences of operations. For the most part, they implement linear algebra computing kernels, which can be effectively parallelized. Hence, a number of customized hardware architectures dedicated to accelerating ML workloads have been proposed in recent years. They exploit reprogrammable logic or fixed-function implementations, such as Tensor Processing Units. A disruptive approach in this space is In-Memory Computing (IMC), for which storage and computation is performed at the same physical location, avoiding the well-known memory wall problem.
@@
Different IMC approaches have been proposed in the literature, operating at varying levels of the memory hierarchy and leveraging different technologies. Among them, Analog In-Memory Computing (AIMC) is emerging as a particularly appealing alternative. A key drawback hampering the widespread adoption of AIMC is its sensitivity to device and circuit-related variations such as temporal conductance variations, device stochasticity, and circuit non-idealities, which can adversely affect accuracy [1]. Many ongoing efforts investigate these effects and propose strategies to mitigate them, including Hardware-Aware (HWA) training [2] and model co-optimization [3].
@@
Addressing the accuracy challenge from a system perspective, we present herein a novel framework that enables the effective exploration of hybrid digital/analog mappings of DL models (Fig. 1), taking into consideration the current limitations of AIMC which affect accuracy. Our framework effectively navigates the trade-off between run-time, performance, and accuracy degradation. Recognizing the exponential relation between the number of MVMs that can be executed in either the digital or analog domain, and the possible mapping solutions, we herein introduce an accuracy-driven heuristic that, by considering the largest MVM operations first (in terms of MAC operations), reduces the exploration problem to that of linear complexity. The main contributions presented in this work are summarized as follows:
@@
  • We present a novel accuracy-driven training framework, able to explore the space of hybrid digital/analog implementations of DNNs. The framework allocates Fully Connected (FC) and Convolutional (CONV) layers to digital or analog resources to minimize run-time while constraining the adverse effect of analog computations on accuracy. Additionally, it is hardware-agnostic, meaning it can be applied to AIMC crossbars adopting different device technologies;
@@
  • We derive a set of optimized hybrid mappings ensuring user-defined accuracy levels for a number of DNNs trained for CIFAR-10 and CIFAR-100 image classification, showing that in all cases a sizable share of their workload can be executed in the analog domain, even for degradation thresholds as low as 0.5% with respect to floating-point precision implementations;
@@
  • We investigate the speedup-accuracy tradeoff achieved by our mappings. To this end, we perform power and performance evaluations on different types of ML workloads using a cycle-accurate simulator based on gem5-X. The considered hybrid system employs a single-core in-order CPU embedding tightly-coupled AIMCs for in-memory acceleration;
@@
  • We recognize and address the real-world system accuracy degradation caused by temporal variation of programmed AIMCs crossbars at a user-desired evaluation time, $t_{eval}$.
@@ II-A In-Memory Computing
Traditional von Neumann computing systems involve separate processing and memory units, which consume significant energy and introduce additional latency when data is moved between them. IMC is a non-von Neumann computational approach, in which computation and storage are performed at the same physical location. From a technology perspective, both SRAM- and DRAM-based IMC architectures have been proposed [1]. Computation in these architectures can be achieved by leveraging charge sharing when concurrently activating multiple memory rows. A promising alternative, which we focus on in this paper, is that of employing Non-Volatile Memory (NVM) devices, organized in a crossbar arrangement [1]. These devices, used to encode network weights in a differential mapping scheme, are employed at the junction of word- and bit-lines, and their resistances are modulated according to the target weight value at compile time.
In AIMC crossbars, inputs are converted to word line voltages using Digital-to-Analog Converters, and currents, representative of MVMs results, are collected across bit lines into Analog-to-Digital Converters. Operations on AIMC crossbars are performed in parallel, and hence, can be realized in constant time complexity. For instance, a 256$\times$256 Phase Change Memory (PCM) crossbar is shown to perform a total of 65,536 MAC operations in 130ns [4]. However, as computation is performed in the analog domain, outputs are affected by non-deterministic noise, deviating from expected results. This is due to a number of device and circuit non-idealities, including temporal and temperature-dependent conductance variations, device stochasticity, ADC mismatch, and IR drop [5].
@@ II-B Hardware-Aware Training
Analog HWA offline training is a common approach to mitigate some of the AIMC device and circuit non-idealities, by making the model more robust for deployment. Additional operations, including weight-noise injection, are added during forward propagation passes. Mathematically, the generally simulated AIMC forward passes can be expressed as
@@
where $f_{\text{adc}}$ and $f_{\text{dac}}$ model the analog-to-digital and digital-to-analog processes, with dynamic scaling and range clipping, and $\xi$ denotes Gaussian noise. $\sigma$ indicates noise sampled from a zero-centered normal distribution. In addition to training noise-resilient weights, it is also possible to train other parameters, such as the input (DAC) bounds of each Word-Line (WL), using back-propagation. We employ the open-source AIHWKIT toolkit [5, 6] to expose these effect in system-level explorations in LionHeart, allowing both the realistic characterization of analog non-idealities and analog HWA retraining. While the toolkit, and by extension, the LionHeart framework can be used to model other device technologies, such as Resistive Random-Access Memory (RRAM), in this paper, we model PCM. Section IV-B3 details the specific assumptions made during analog HWA retraining.
II-C Related Work
@@
Related works, which perform application mapping on AIMC- accelerated systems, consider resource limitations as their main (or only) constraint. PUMA [7] employs compiler analysis passes to define which parts of the application have the most potential for speedup, while AERO [8] formulates this same problem as a cost function minimization. Potential issues introduced by non-idealities and noisy environments, characterizing AIMC devices, are overlooked in the above-mentioned works. They are similarly disregarded in [9] and [10], where all FC and CONV layers are executed using AIMC tiles, without attempting to control induced accuracy degradation. Some consideration of the effect of analog computing on accuracy is included in [11], where the authors propose a hybrid solution where the first and last layers of DNNs are executed on digital resources, postulating that these two layers are more sensitive to noise. Two recent works present input channel-wise weight selection methods to map to heterogeneous analog-digital resources. The first, HybridAC [12], does so without employing any re-training strategy. The input is assumed to already be HWA trained, and some layers are confined to be implemented using digital resources. The second, ODiMO [13], relies on the unique approach of training a supernetwork. Analog computation is modelled using 2-bit quantization. However, both of these methods do not consider (i) the variability of analog tiles as well as (ii) the temporal evolution of analog weights.
@@
The complexity of devising high-performance, but accuracy-constrained mappings, is compounded by the varying size and sensitivity towards noise of the different layers in a DNN structure. Moreover, the number of possible mappings makes exhaustive exploration impractical, even for simple cases, as discussed, in the context of hardware quantization, in [14]. For example, even when layer-wise mapping is considered, a relatively low-depth model like AlexNet, composed of just seven layers, would therefore present $2^{7}$ possible analog-digital combinations. These would all have to be trained and tested to arrive at an exact solution to the mapping problem. In contrast to the aforementioned works, we employ a heuristic-based layer-wise mapping strategy which scales linearly with respect the number of layers, as it only consider one target candidate layer at a time.
@@ III LionHeart framework
Model Weights Top-1 Acc. (%)
(# Mappable Layers)
CIFAR-10
CIFAR-100
CIFAR-10
CIFAR-100
@@ IV-A Experimental Setup
The CIFAR-10 and CIFAR-100 datasets were chosen for evaluation. We collect results using these datasets for the five DNN architectures targeting edge devices as listed in Table I, where the size of the last FC layer is modified according to the number of output classes (10 and 100, respectively). The number of weights of each DNN ranges from 77,968 to 15,237,698, and hence contains both under- and over-parameterized models for the tasks. Data is split into training, validation, and testing sets. The validation dataset was obtained by splitting the original training dataset into two separate datasets, comprising 90% of training images and the remaining 10% of validation images. The same randomly sampled validation set was used for all experiments.
Figure 3: (a-e) The MAC ratio (as a percentage) for all configurations between the number of MACs performed using AIMC and the total number of MACs. (f-j) The corresponding test set accuracies at $t=$1d.

### v3

@@ Abstract
When arranged in a crossbar configuration, resistive memory devices can be used to execute Matrix-Vector Multiplications, the most dominant operation of many Machine Learning (ML) algorithms, in constant time complexity. Nonetheless, when performing computations in the analog domain, novel challenges are introduced in terms of arithmetic precision and stochasticity, due to non-ideal circuit and device behaviour. Moreover, these non-idealities have a temporal dimension, resulting in a degrading application accuracy over time. Facing these challenges, we propose a novel framework, named LionHeart, to obtain hybrid analog-digital mappings to execute Deep Learning (DL) inference workloads using heterogeneous accelerators. The accuracy-constrained mappings derived by LionHeart showcase, across different Convolutional Neural Networks (CNNs) and one transformer-based network, high accuracy and potential for speedup. The results of the full system simulations highlight run-time reductions and energy efficiency gains that exceed 6$\times$, with a user-defined accuracy threshold for a fully digital floating point implementation. LionHeart is open-sourced here: https://github.com/IBM/lionheart.
@@ I Introduction
Deep Learning (DL)-based solutions have been applied in various industrial and scientific applications, including computer vision, natural language processing, and speech recognition. Nowadays, State-Of-The-Art (SOTA) DL models require significant memory to store their parameters (i.e., weights and biases), and billions of Multiply-Accumulate (MAC) operations per inference. These extreme computing and memory requirements pose a challenge to deploy DL workloads on edge devices, whose form factor and energy budget require limited computing capabilities and memory size. However, the execution of DL workloads at the edge represents a very attractive prospect, as it would allow the processing of data samples close to where it is collected, resulting in several benefits, including reduced latency, a higher degree of security and privacy, improved bandwidth efficiency, as well as increased resiliency. To enable DL within the tiny resource envelopes of edge devices, workload optimization is key. To this end, many research works have proposed dedicated hardware solutions, customized for DL workloads [1, 2].
Figure 1: LionHeart heterogeneously maps layers of ML networks to digital or analog resources, maximizes performance by exploiting AIMC acceleration, while at the same time abiding to accuracy constraints.
@@
Such accelerators leverage the structured organization of ML algorithms, such as Deep Neural Networks and transformers, which are composed of layers sequences that, for the most part, implement linear algebra computing kernels. In turn, these can be effectively supported by hardware architectures by exploring their parallelism, whether employing reprogrammable logic or fixed-function implementations, such as Tensor Processing Units [3]. A disruptive approach in this space is In-Memory Computing (IMC), for which storage and computation is performed at the same physical location, avoiding the well-known memory wall problem.
Different IMC approaches have been proposed in the literature, operating at varying levels of the memory hierarchy and leveraging different technologies. We summarize this landscape in Section II. Among them, Analog In-Memory Computing (AIMC) is emerging as a particularly appealing alternative. The AIMC computing paradigm is based on crossbar architectures, with devices having variable conductance connecting rows and columns. By leveraging Ohm’s law and Kirchhoff’s current law, AIMC accelerators can be used to perform matrix-vector operations between an input vector (encoded using Word-Line (WL) voltages) and a matrix (encoded using conductance) in linear time [4], i.e., the time required to write inputs and read-back outputs.
@@
An important drawback hampering the widespread adoption of AIMC is its sensitivity to device and circuit-related variations such as temporal conductance variations, device stochasticity, and circuit non-idealities, which can adversely affect accuracy [2]. Many ongoing efforts investigate these effects and propose strategies to mitigate them, including Hardware-Aware (HWA) retraining [5] and model co-optimization [6]. However, none of these works adequately addresses the accuracy challenge from a system perspective, i.e. how AIMC acceleration can be leveraged to speed up ML execution while faithfully modeling and controlling the ensuing noise-induced accuracy degradation.
@@
We herein aim at filling this gap. To this end, we present a novel framework that enables the effective exploration of hybrid digital/analog mappings of DL models (Fig. 1), taking into account the current limitations of AIMC which affect accuracy. Our framework effectively navigates the trade-off between run-time, performance, and accuracy degradation. Recognizing the exponential relation between the number of MVMs that can be executed in either the digital or analog domain, and the possible mapping solutions, we herein introduce an accuracy-driven heuristic that, by considering the largest MVM operations first (in terms of MAC operations), reduces the exploration problem to that of linear complexity. The main contributions presented in this work are summarized as follows:
@@
  • We present a novel accuracy-driven training framework, able to explore the space of hybrid digital/analog implementations of DNNs. The framework allocates different layers to digital or analog resources to minimize run-time while constraining the adverse effect of analog computations on accuracy. Additionally, it is hardware-agnostic, meaning it can be applied to AIMC crossbars adopting different device technologies;
  • We derive a set of optimized hybrid mappings ensuring user-defined accuracy levels for a number of DNNs trained for CIFAR-10 and CIFAR-100 image classification, in addition to a transformer-based network for Stanford Question Answering Dataset (SQuAD) – a Natural Language Processing (NLP) task;
@@
  • We showcase, that for all of these networks, a sizable share of their workload can be executed in the analog domain, even for degradation thresholds as low as 0.5% with respect to floating-point precision implementations;
@@
  • We investigate the speedup-accuracy tradeoff achieved by our mappings. To this end, we perform power and performance evaluations on different types of ML workloads using a cycle-accurate simulator based on gem5-X [4];
  • We compare our optimal mappings to those obtained from prior work (DIANA [7] and Harmonica [8]), and evaluate the variability of the achieved mappings, in addition to their accuracy and MAC ratio;
  • For one network, we hardware-aware train all possible mappings and establish a upper-bound for comparison;
@@
  • Finally, we recognize and address the real-world system accuracy degradation caused by temporal variation of programmed AIMCs crossbars at a user-desired evaluation time, $t_{eval}$.
@@ II-A In-Memory Computing
Traditional von Neumann computing systems involve separate processing and memory units, which consume significant energy and introduce additional latency when data is moved between them. IMC is a non-von Neumann computational approach, in which computation and storage are performed at the same physical location. Several different IMC architectures, which aim to blend computation and storage, have emerged in recent years [9]. Architectures can utilize IMC at different levels of the memory hierarchy, for example, when interfacing main memory [10], as part of smart caches [11], or as functional units that augment processor pipelines and register files [4]. While we consider this last scenario in our experimental evaluation when evaluating speedups, our methodology is agnostic with respect to the integration choice, because we abstract architectural aspects when performing application mapping to analog/digital resource. From a technology perspective, both SRAM- and DRAM-based IMC architectures have been proposed [2]. Computation in these architectures can be achieved by taking advantage of charge sharing when simultaneously activating multiple memory rows. A promising alternative, which we focus on in this paper, is that of employing Non-Volatile Memory (NVM) devices, organized in a crossbar arrangement [2]. These devices, used to encode network weights in a differential mapping scheme, are employed at the junction of word- and bit-lines, and their resistances are modulated according to the target weight value at compile time.
As depicted in Fig. 2, AIMC crossbars encode inputs as word line voltages using Digital-to-Analog Converters, and currents, representative of MVMs results, are collected across bit lines into Analog-to-Digital Converters. Operations on AIMC crossbars are performed in parallel, and hence, can be realized in constant time complexity. For instance, a 256$\times$256 Phase Change Memory (PCM) crossbar is shown to perform a total of 65,536 MAC operations in 130ns  in Le Gallo et al.[12]. However, as computation is performed in the analog domain, outputs are affected by non-deterministic noise, deviating from expected results. This is due to a number of device and circuit non-idealities, including temporal and temperature-dependent conductance variations, device stochasticity, ADC mismatch, and IR drop [5].
Figure 2: (a) Depiction of an AIMC tile and its underlying compute mechanism. Each memristor, as depicted, is representative of a unit cell [^2] AIMC tiles may be realized using different architectures [13, 14], most prominently, pseudo-crossbar or conventional memory, where the input injection ways are different, and the single-bit-multi-bit encoding capabilities are also different. LionHeart is architecture agnostic..
Figure 3: Mapping of (a) FC and (b) CONV layers to device conductances. Weights are linearly scaled and mapped between $G_{min}$ and $G_{max}$. (c) Mapping/execution flow of a self-attention block when the maximum analog MAC ratio is achieved.
@@ II-B Hardware-Aware Training
Analog HWA offline training is a common approach to mitigate the AIMC device and circuit non-idealities, by making the model more robust for deployment. When performing HWA training, weight-noise is injected during forward propagation passes. Mathematically, this is expressed as follows:
@@
where $f_{\text{adc}}$ and $f_{\text{dac}}$ model the analog-to-digital and digital-to-analog processes, with dynamic scaling and range clipping, and $\xi$ denotes Gaussian noise. $\sigma$ indicates noise sampled from a zero-centered normal distribution. In addition to training noise-resilient weights, it is also possible to train other parameters, such as the input (DAC) bounds of each tile, using back-propagation. We employ the open-source AIHWKIT toolkit [5] to expose these behaviours using system-level explorations in LionHeart, allowing both the realistic characterization of analog non-idealities and analog HWA retraining. While the toolkit, and by extension, the LionHeart framework can be used to model other device technologies, such as Resistive Random-Access Memory (RRAM), in this paper, we model PCM. The PCM device technology was chosen as compared to other device technologies, it utilizes fab–friendly materials and can achieve high conductance precision. Section IV-B3 details the specific assumptions made during analog HWA retraining.
II-C Computational Complexity and Execution of Convolutional and FC Layers Using IMC
FC layers execute a number of MAC operations which is linearly proportional to the number of input and output elements. Specifically, they have a complexity of $\mathcal{O}(MN)$, with $M$ and $N$ representing the number of input and output features, respectively. A CONV layer that applies $F$ filters to compute an output feature map of size $W_{o}\times{}H_{o}\times{}F$ requires a number of MAC operations in the order of $\mathcal{O}(W_{o}H_{o}CK^{2}F)$, where $K$ represents the kernel size and $C$ the number of input channels.
Weights of FC layers can be computed directly using the MVM operation, whereas weights of CONV layers can be transformed to MVMs using the im2col algorithm [15] (see Fig. 3). Using im2col, weights of CONV layers are unfolded and indexed as a matrix, where each column of the matrix contains the unrolled values of a filter. For both layer types, unfolded weight matrices are typically much larger than the attainable dimensions of AIMC tiles. For both layer types, unfolded weight matrices are typically much larger than the attainable dimensions of AIMC tiles, e.g., all convolutional layers of ResNet20 require 32 256$\times$256 tiles. Hence, tiling can be employed to partition large matrices into a number of discrete tiles. This strategy requires an additional computing step to aggregate partial results, which is performed per tile, rather than per element [4]. Unless otherwise stated, we utilize a distributed weight mapping scheme, in which larger matrix are evenly distributed among tiles.
For self-attention blocks, we consider mapping the first three FC layers of each head and the last FC layer after concatenation. FC layers within the scaled-dot product attention are not considered, as they have dynamic weights, which for AIMC require constant reprogramming during infererence.
II-D Related Work
@@
Most related work, which performs application mapping on AIMC-accelerated systems, consider resource limitations as their main (or only) constraint. PUMA [16] employs compiler analysis passes to define which parts of the application have the most potential for speedup, while AERO [17] formulates this same problem as a cost function minimization. Potential issues introduced by non-idealities and noisy environments, characterizing AIMC devices, are overlooked in the above-mentioned works. They are similarly disregarded in [18] and [19], where all FC and CONV layers are executed using AIMC tiles, without attempting to control induced accuracy degradation. Some consideration of the effect of analog computing on accuracy is included in [7], where the authors propose a hybrid solution where the first and last layers of DNNs are executed on digital resources, postulating that these two layers are more sensitive to noise.
Two recent works present input channel-wise weight selection methods to map to heterogeneous analog-digital resources. The first, Harmonica [20], does so without employing any re-training strategy. The input is assumed to already be HWA trained, and some layers are confined to be implemented using digital resources. For channels of unconfined layers, the Hessian sensitivity [21] is determined. These are then ordered from the most sensitive to the least sensitive. The (next) most sensitive channel is converted to digital and the accuracy of the network is determined. This process is repeated until the accuracy of the network drops below a pre-defined threshold. Harmonica is described in full in Algorithm 1 of [20].
The second, ODiMO [22], relies on the unique approach of training a supernetwork, where for each channel, both digital and AIMC implementations are trained simultaneously. After the network is trained, for each channel, either the digital or AIMC implementation is pruned. Analog computation is modelled using 2-bit quantization. Both of these methods do not consider (i) the variability of analog tiles as well as (ii) the temporal evolution of analog weights (conductances).
@@
The complexity of devising high-performance, but accuracy-constrained mappings, is compounded by the varying size and sensitivity towards noise of the different layers in a DNN structure. Moreover, the number of possible mappings makes exhaustive exploration impractical, even for simple cases, as discussed, in the context of hardware quantization, in [23]. For example, even when layer-wise mapping is considered, a relatively low-depth model like AlexNet, composed of just seven layers, would therefore present $2^{7}$ possible analog-digital combinations. All of these would have to be trained and tested to arrive at an exact solution to the mapping problem. In contrast to Harmonica and ODiMO, we employ a heuristic-based layer-wise mapping strategy, which scales linearly with respect to the number of layers, as it only considers one target candidate layer at a time.
@@ III LionHeart framework
Model Weights Top-1 Acc./F1 (%)
(# Mappable Layers) CIFAR-10 CIFAR-100 CIFAR-10 CIFAR-100
@@
While not investigated in this paper, it is possible to extend the methodology of LionHeart to co-apply different techniques to further improve accuracy, e.g., tunable ADC resolutions in RAELLA [25].
@@ IV-A Experimental Setup
For the considered DNNs, similarly to [7] and [8] we evaluate the performance of LionHeart on the CIFAR-10 dataset [26]. Moreover, we also provide outcomes on the more challenging CIFAR-100 dataset. [26]. We consider five DNN architectures targeting edge devices, as listed in Table I, where the size of the last FC layer is modified according to the number of output classes (10 and 100, respectively).
While the degree to which a DNN architecture is parameterized for a given dataset is difficult to determine quantitatively, the DNN architectures were selected such that they range from being under- to over-parameterized so that they are representative of a broader range of networks and chosen tasks. To demonstrate the generality of LionHeart beyond CNNs, we also apply it to the MobileBERT transformer benchmark, processing on the SQuAD dataset [27] [^3] We refer to the SQuAD 1.1 dataset as SQuAD throughout the paper..
We collect results using CIFAR-10 and CIFAR-100 for the first five DNN architectures targeting edge devices as listed in Table I. Additionally, we collect results using SQuAD for MobileBERT. The number of weights of these networks range from $\approx$78K to $\approx$25M. Data is split into training, validation, and testing sets. For CIFAR-10 and CIFAR-100, the validation dataset was obtained by splitting the original training dataset into two separate datasets, comprising 90% of training images and the remaining 10% of validation images. The same randomly sampled validation set was used for all experiments. For SQuAD, the pre-existing development set (comprising 10% of all inputs) was used as the validation set. The test sets of all datasets were solely used for final model evaluation.
Figure 5: For (a,b) ResNet8, (c,d) ResNet20, (e,f) AlexNet, (g,h) VGG16, (i,j) MobileNetV2, and (k,l) MobileBERT, the MAC ratio (as a percentage) and corresponding test set accuracies (F1 scores for SQuAD) at $t=$1d. The MAC ratio quantity represents the ratio between the number of MACs performed using AIMC and the total number of MACs
TABLE II: Shared training hyper-parameters for the CIFAR-10/ CIFAR-100 and SQuAD tasks.
Parameter CIFAR-10/CIFAR-100 SQuAD
Optimizer SGD + Momentum
Loss Function Cross Entropy
LR Scheduler CosineAnneling Linear
T_max 50 3
Batch Size 256 8
Digital
LR 0.057 0.05
Momentum 0.867 0.99
Weight Decay 0 0.00054
Analog
LR 0.024 0.02
Momentum 0.775 0.9
Weight Decay 0 0.003